OpenStudy (anonymous):

what is the area, in square millieters, of a right triangle with a hypotenuse of 68 millimeters and leg of 60 millimeters? a. 2,040 mm^2 b. 1,920 mm^2 c. 1,088 mm^2 d. 960 mm^2

OpenStudy (cwrw238):

find length of the other leg using pythagoras = sqrt(68^2 - 60^2) then the area = 0.5 * 60 * this value

OpenStudy (anonymous):

30...?

OpenStudy (cwrw238):

what - the other leg - pretty close

OpenStudy (cwrw238):

so area = 30 * length of other leg

OpenStudy (cwrw238):

sqrt (1024)

OpenStudy (cwrw238):

- thats the other leg

OpenStudy (anonymous):

ok